Question i (i point) solve and check:-4 log2 (2-3) 4 . [Hin a) x=2; check-4(-1)-4 b) x=-1; check 1082 (4 + 12) c) x=5; checklog2 24 4 = 4 d) x=3.5; check-4(-1)-4 Question 2 (1 point)
Answer #1

Given -4log2(x-3)=4

Then log2(x-3)=-1

x-3=2-1

x-3=1/2

x-3=0.5

Therefore x =3.5

Check -4log 2(3.5-3)= -4log22-1 =-4(-1) =4

Therefore answer is option d) x=3.5 ,check -4(-1)=4